A=50DEGREE,B=30DEG, c=9
solve each triangle described below. round measures of sides and angles to the nearest tenth.

Answer by stanbon(75887)

A=50DEGREE,B=30DEG, c=9
solve each triangle described below. round measures of sides and angles to the nearest tenth.
--------------------
C = 180-(50+30) = 100
Use the law of sines:
a/sin(50) = 9/sin(100)
a = 9sin(50)/sin(100)
a = 4.6634
------------
b/sin(30) = 9/sin(100)
b = 9sin(30)/sin100
b = 17.5610
================
Cheers,
Stan H.